OpenStudy (anonymous):

A car travels 60.0 miles per hour. What is its velocity in m/s? (1 km = 0.621 mi) 0.01 m/s 0.03 m/s 9.66 m/s 26.8 m/s

OpenStudy (anonymous):

this is the easiest way to convert one unit into another: writing the same thing in 2 different units such that the unit that you don't want get's cancelled. Here's what i mean: \[\text{60 mile an hour = }60\frac{mile}{hour}*\frac{1.61km}{1 mile}*\frac{1000m}{1km}*\frac{1hour}{3600s}\\\\\text{Now you cancel the mile in the numerator & denominator. Same with that km and hour.}\\\\\text{what you get is } \text{60 mile an hour = }\frac{60*1.61*1000}{3600}\frac{m}{s}\] Simplify to get your answer.